OpenStudy (anonymous):

LAST ONE I PROMISE!!! 179.2 is 32% of what number? FREE MEDALS

OpenStudy (anonymous):

(179.2/x)*100=32 179.2/x=0.32 179.2=0.32x 179.2/0.32=x x=560

OpenStudy (anonymous):

I officially LOVE you:)

OpenStudy (anonymous):

Lolz.

OpenStudy (anonymous):

I have a girl already